OpenStudy (anonymous):

Can anyone help me with this one please simplify 3/5+(-1/7) all over 1and1/2

OpenStudy (phi):

\[\frac{\frac{3}{5}-\frac{1}{7}}{1\frac{1}{2}}\] ?

OpenStudy (phi):

if so, make the denominator an improper fraction \[ 1 \frac{1}{2}= \frac{3}{2} \] multiply top and bottom by 2/3 to get rid of the denominator \[\frac{(\frac{3}{5}-\frac{1}{7})\cdot \frac{2}{3}}{\cancel{\frac{3}{2}}\cdot\cancel{\frac{2}{3}}}\]

OpenStudy (phi):

Can you finish this?

OpenStudy (anonymous):

I'll give it a try

OpenStudy (phi):

multiplying fractions http://www.khanacademy.org/video/multiplying-fractions?topic=fractions

OpenStudy (anonymous):

If I did this correctly the answer would be 32/105 ?

OpenStudy (anonymous):

Thanks for the links

OpenStudy (phi):

\[(\frac{3}{5}-\frac{1}{7})= \frac{3}{5}\cdot\frac{7}{7}-\frac{1}{7}\cdot \frac{5}{5}= \frac{21}{35}-\frac{5}{35}\]

OpenStudy (phi):

which becomes 16/35 times 2/3 = 32/105

OpenStudy (phi):

Looks like you know what you are doing.